OpenStudy (anonymous):

What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x

OpenStudy (anonymous):

Assuming p = period of the function or 2pi, sin(x+p) = sin(x). Thus, answer B, or sinx, would be the correct answer.

OpenStudy (anonymous):

alright but how would I get that

OpenStudy (jdoe0001):

I gather he meant to say "pi", for \(\large \pi \)

OpenStudy (anonymous):

ohhhh okay thanks man/women.

OpenStudy (jdoe0001):

so, use the sum trig identity -> http://www.sosmath.com/trig/addform/img1.gif for sine

OpenStudy (jdoe0001):

and see what you get

OpenStudy (anonymous):

but how would I know what to do if the only information there is sin(x+p)

OpenStudy (jdoe0001):

well, off the trig identities there, which one do you think we could use?

OpenStudy (anonymous):

i am terribly sorry but i am in james Madison high school from ashworth

OpenStudy (jdoe0001):

you're sorry you're at that HS.... or you're sorry for James Madison who's been dead for about 190 years?

OpenStudy (anonymous):

sorry for being at this high school the teachers don't really teach any thing. but you would use the sin(a+b)

OpenStudy (jdoe0001):

right, as you can see in the picture, the identity is sin(a + b) = sin(a)cos(b)+cos(a)sin(b) thus \(\bf sin(x+\pi)\\\quad \\ \quad \\ sin(x+\pi) \implies sin(x)cos(\pi)+cos(x)sin(\pi)\\ \quad \\\quad \\ \textit{what's the }sin(\pi) \textit{ and the } cos(\pi)\quad ?\)

OpenStudy (jdoe0001):

if you check your Unit Circle, you'll see what those two are

OpenStudy (anonymous):

what unit circle?

OpenStudy (jdoe0001):

hehh, man, that'.s... anyhow... if you need a Unit Circle, as you should have had one, many online, try this one --> http://i.stack.imgur.com/r8uHr.gif

OpenStudy (jdoe0001):

so.. \(\bf \textit{what's the }sin(\pi) \textit{ and the } cos(\pi)\quad ?\)

OpenStudy (anonymous):

dude i am totally lost by looking at this but i would have to guess it is 180 for both?

OpenStudy (jdoe0001):

heheh, that Unit Circle, or any good unit circle, will have the degrees, 180 yes, as well as the Radians units, notice it shows as \(\large 180^o = \pi\)

OpenStudy (anonymous):

yes. i do

OpenStudy (anonymous):

so would you plug 180 into sin(pi) and cos(pi)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

alright so i got -1

OpenStudy (jdoe0001):

-1? for...?

OpenStudy (anonymous):

for sin and cos

OpenStudy (jdoe0001):

hmm.... on the Unit Circle, there's a pair of values, just like (x, y) coordinates the pair is (x, y) also or (cosine, sine) pair so one value is for cosine, the other is for sine

OpenStudy (jdoe0001):

if you notice the top-right-hand corner, it shows the (cos, sin) pair reference

OpenStudy (anonymous):

oh oh okay now i see it so it would be (-1,0)

OpenStudy (jdoe0001):

that means that \(\bf \large cos(\pi) = -1 \qquad sin(\pi) = 0\) so replace that in \(\bf sin(x)cos(\pi)+cos(x)sin(\pi) \) see what you get

OpenStudy (anonymous):

so you would get sinx

OpenStudy (jdoe0001):

\(\bf sin(x)cos(\pi)+cos(x)sin(\pi) \implies sin(x)-1+cos(x)0\\\quad \\ sin(x)-1+cos(x)0 \implies sin(x)-1+0 \implies -sin(x)\) anything * 0 = 0

OpenStudy (anonymous):

alright now this all makes sense

OpenStudy (anonymous):

thanks man appreciate it

OpenStudy (jdoe0001):

yw